if I had to write out thirty six squared divided by six cubed would I write it like 36^2/6^3 ot use the divide symbol
Answer by ptaylor(2198)

Looks good to me!
Now if you wanted to solve it, you could write (36*36)/(6*6*6)=
(36*36)/(36*6)=36/6=6
Hope this helps---ptaylor